Keaton Inc. has come out with a new and improved product, and is expected to have an EPS of $4.2 and an ROE of 20%. It will maintain a plowback ratio of 35%. Investors expect a 12% rate of return on the stock. Assuming Keaton's current value is measured with the constant growth DDM, compute the present value of growth opportunities for Keaton.

*Round your answer to TWO decimal places.

From the earlier problem with Keaton, suppose the present value of growth opportunities = 0. If everything else remains constant, find the cost of equity for Keaton and explain why.
Answer #1

Growth Rate = 0.35(0.20) = 7.00%

Interest Rate = 12%

D1 = (1 - 0.35)(4.20) = $2.73

Using Constant Growth Model,

Stock Price = 2.73/(0.12 - 0.07) = $54.60

PVGO = Stock Price - E1/r

PVGO = 54.60 - 4.20/0.12

PVGO = $19.60